# On the Hyperparameter Loss Landscapes of Machine Learning Algorithms

### v1

@@ 4.3 HPO Landscapes with Different Fidelities
Figure 2 (c) shows the low-fidelity test loss landscapes (denoted as $\mathcal{L}_{\text{test}_{\text{LF}}}$) for each model (using $10$ epochs for FCNet, and $10\%$ training data for others). From the plots, we could see that $\mathcal{L}_{\text{test}_{\text{LF}}}$ landscapes are highly consistent with $\mathcal{L}_{\text{test}}$ in terms of both structural characteristics and performance rankings. More specifically, as reflected in Figure 3, all measured FLA metrics of $\mathcal{L}_{\text{test}_{\text{LF}}}$ landscapes showed little difference compared to $\mathcal{L}_{\text{test}}$ landscapes across all studied scenarios. For performance rankings, Figure 4 (b) depicts the distribution of the $3$ similarity indicators between $\mathcal{L}_{\text{test}_{\text{LF}}}$ and $\mathcal{L}_{\text{test}}$ across all datasets for each model. We could observe a high Spearman correlation (median $>0.85$) between $\mathcal{L}_{\text{test}}$ and $\mathcal{L}_{\text{test}_{\text{LF}}}$ for all models, and the $\gamma$-set similarities between the top-$10\%$ configurations are also prominent, with medians larger than $60\%$. These imply that $R(\mathcal{L}_{\text{test}})$ and $R(\mathcal{L}_{\text{test}_{\text{LF}}})$ are highly consistent for the majority of our studied scenarios and there is large overlap between the promising regions of the two landscapes. In addition, the Shake-up scores yield low values ($\text{median}<0.1$), suggesting that on average the difference between $R(\mathcal{L}_{\text{test}})$ and $R(\mathcal{L}_{\text{test}_{\text{LF}}})$ is less than $10\%$. Additional results on FCNet, NAS-Bench-101 in Appendix D and Appendix C respectively are also consistent with our findings here.
4.4 HPO Landscapes Across Datasets
@@
Figure 6: Importance of top-$5$ HPs as determined by functional ANOVA method for each model.
In addition to performance rankings, Figure 6 illustrates the contribution of each HP and their interactions to model performance assessed by the functional ANOVA method. The results indicate that some (combination of) HPs are typically important for many datasets for a given model. For example, learning rate consistently contributes a large portion of variance to model performance for LightGBM, and its interactions with the number of leaves and estimators are also important. These observations are similar with van Rijn & Hutter (2018), which find also conclude that certain HPs of a ML model are important for a wide spectrum of datasets by analyzing meta-data from OpenML platform.
As discussed in Section 4.1, HP loss landscapes often involve a large number of none-improvement moves, especially near the optimum. We also see that there is clear division between the promising regions and the poorly-performaning ones. Therefore, leveraging prior knowledge from previously tasks should be able to greatly expedite the searching process by means like warm-starting HPO from good configurations or carefully selecting candidate HPs and crafting the search space. More importantly, based on our results, we note that this should not be only limited to similar tasks defined under certain rules, since they may not always available. On the other hand, seemingly different tasks could still provide informative information as we see above. Our additional results for FCNet on $4$ datasets, and NAS-Bench-201 across CIFAR-10/100 as well as ImageNet datasets (Appendix C), also revealed similar highly-transferable conclusions.
5 Disuccsions and Conclusions
By conducting large-scale exploratory analysis on $1,500$ HP landscapes of $6$ ML models with over $11$M model configurations under different scenarios, this paper reveals an unified portrait of their topographies in terms of smoothness, neutrality and modality. We also show that these properties are highly transferable across datasets and fidelities, and thus provide fundamental evidence to support the effectiveness of transfer and multi-fidelity methods, which in privious practices, is mainly based on intuition. However, while our findings are observed for the majority studied scenarios, we do observe some exceptions. For example, most landscapes inspected reveal a nearly unimodal structure, but some of them can have dozens to a few hundreds of local optima with non-negligible basin sizes (e.g., FCNet). Also, there are cases when landscapes under lower fidelities or on a different task reveal very different patterns, as shown by the long tails of the similarity distributions in Figure 4. Further explorations by interrogating the relationship between dataset characteristics may provide a even more comprehensive understanding of the HPO landscape.
Our developed FLA framework in this work, has shown great potential for enabling both qualitative and quantitative understanding for a wider range of AutoML problems. While it currently relies on large-scale, pre-evaluated data points for landscape construction, we think it can be a promising direction to integrate it with existing AutoML frameworks, and thus allow for on-the-fly analysis of problem landscapes, thereby allowing it to be accessible by a broader range of stakeholders.
Appendix A Background and Related Work
Fitness Landscape of BBOPs. A central topic in optimization concerns how the intricate interplay of decision variables determine the targeted objective values. For example, geneticists and biologists are interested in the relationship between genotypes and the phenotypes and functions of organisms that may affect their evolutionary success (De Visser & Krug, 2014). In chemical physics, scientists seek to explore how the energy of a system is dependent on its various configurational or conformational states (Brooks III et al., 2001). The fitness landscape can be envisioned as a surface over the high-dimensional space as formed by decision variables, where the objective values are represented as the elevation of the landscape. The structure of this surface describes the spectrum of possible objectives values across the variable space and thus strongly influences the optimization. In addition, searching trajectories of problem solvers can be thought as strategic walks on the corresponding problem landscape. Successful applications of this metaphor to analyze black-box systems have advanced the understanding of many fields (He & Liu, 2016; Puchta et al., 2016; Shires & Pickard, 2021; Brooks III et al., 2001). Most related to our work is the fitness landscape analysis (FLA) for black-box optimization problems (BBOPs) (Malan, 2021). Many metrics that quantify the structural characteristics of fitness landscapes have been developed over the years to describe the topography of BBOP landscapes, e.g., autocorrelation, neutrality, modality, epistasis variance, evolvability, etc.
Fitness Landscape on other ML Systems. In addition to the works that focus on analyzing HP and AutoML landscapes introduced in Section 1, there are also works which applied FLA to study other ML systems, including the neural architecture search (NAS) spaces (Rodrigues et al., 2020; Traoré et al., 2021) and the HP landscapes of reinforcement learning (RL) (Eimer et al., 2023; Mohan et al., 2023). We note that these works are somewhat orthogonal to ours, as we focus on the HP landscapes of ML algorithms. In addition, there is also another line of work on investigating the loss landscapes of neural networks (Rakitianskaia et al., 2016; Rodrigues et al., 2022; van Aardt et al., 2017). These works, though seemingly similar, are focused on the neural network losses which are the objectives of weight optimizers during model training, and are thus significantly different from ours.
HPO Landscape Visualization. The ability to visualize HP loss landscapes is crucial for enabling intuitive understandings of their complex topologies, which is however, far from trivial in practice due to their high-dimensional nature. Current study methods usually bypass this issue by exmaining only two HPs each time (see, e.g., Akiba et al. (2019); Ørebæk & Geitle (2021)). However, such techqniues can not provide an unified illustration of the total interactive effects of all interested HPs. In the evolutionary computation, there are have been successful attempts on applying dimensionality reduction techniques to visualize the fitness landscape of BBOPs (Michalak, 2019; Huang & Li, 2023; Sass et al., 2022; Walter et al., 2022). Similar work has also been done in the domain of chemical physics (Shires & Pickard, 2021), where UMAP and $t$-SNE are used to visualize the so-called energy landscape.
Overfitting in Machine Learning. To empirically investigate the presence of overfitting in real-world ML applications, Roelofs et al. (2019) analyzed massive submission meta-data on the Kaggle platform, and their results suggest little evidence of overfitting due to testset reuse. However, since the Kaggle data is constituted by a diverse set of ML models with very different HP configurations that are not comparable with each other, it prohibits further analysis on which HP configurations are more likely to lead to overfitting.
Hyperparameter Importance and Interaction. Understanding which HPs influence model performance to what extend can provide valuable insights into the tuning strategy (Probst et al., 2019). To analyze importance of hyperparameters, one could either use models that are inherently interpretable, e.g., decision trees (Quinlan, 1986), or apply model-agnostic methods such as functional ANOVA (Hutter et al., 2014a; Watanabe et al., 2023b). Built on this technique, van Rijn & Hutter (2017) compared the performance of a variety of algorithms on large set of OpenML datasets and stated that the same hyperparameters were typically important across datasets.
Appendix B Details of the Landscape Analysis Framework
Figure 7: High-level conceptual workflow of our proposed HP loss landscape analysis framework.
Figure 7 illustrates the conceptual workflow of our proposed HP loss landscapes analysis framework in Section 2. From a high level, our landscape visualizations (Appendix B.1) seek to provide a general sketch of the landscape topography. The observed patterns can then be quantitatively verified via a set of dedicated FLA metrics characterizing complementary aspects of landscape properties (Appendix B.2), while landscapes visualizations, at the same time, can assist the intuitive interpretation of the numerical figures. In addition to these, we also apply $3$ dedicated similarity measures to quantify the consistency of configuration performance across landscapes (Appendix B.4).
@@ B.1 Landscape Visualization Method
HOPE Node Embedding. To preserve the intrinsic neighborhood relationship of HP configurations, our proposed landscape visualization method first needs to learn a low-dimensional embedding for each configuration in the landscape (i.e., each node in the graph). In this paper, we choose HOPE [^1] We use the implementation in Karateclub package. node embedding method to serve this purpose, as it is able to capture asymmetric high-order proximity in directed networks. Specifically, in a directed network [^2] Most manipulations and analyses of HP loss landscapes as complex networks are grounded on NetworkX and Pandas package., if there is a directed link from vertex $v_{i}$ to vertex $v_{j}$ and from vertex $v_{j}$ to vertex $v_{k}$, it is more likely to have a link from $v_{i}$ to $v_{k}$, but not from $v_{k}$ to $v_{i}$. In order to preserve such asymmetric transitivity, HOPE learns two vertex embedding vectors $U^{s},U^{t}\in\mathbf{R}^{|V|\times d}$, which is called source and target embedding vectors, respectively. After constructing the high-order proximity matrix $S$ from $4s$ proximity measures, i.e., Katz Index, Rooted PageRank, Common Neighbors and AdamicAdar. HOPE finally learns vertex embeddings by solving the a matrix factorization problem:
@@
UMAP Dimensionality Reduction. While HOPE (as well as any other node embedding methods) could generate vectorized embeddings for configurations, these embeddings typically are still high-dimensional and not compatible for $2$D visualization. To cope with this, we further apply UMAP [^3] We use the implementation in UMAP package. to project the HOPE embeddings into a $2$D space, which is based on three assumptions, and here we provide brief discussion on whether they hold for our case:
  • Data are uniformly distributed on Riemannian manifold. As the distribution of HOPE embeddings essentially depends on the connectivity pattern of the graph, for HP loss landscapes, we can expect this assumption to hold. This is because the number of neighbors for each node in the graph would largely stay the same based on our neighborhood definition. Therefore, there will be no significance variation of density across the graph, and the HOPE embeddings in turn should approximately have a uniform distribution.
  • The Riemannian metric is locally constant. The distances between HOPE embeddings correlate directly with the local connectivity patterns within the network. Given the stability in neighborhood structures as we discussed above, it is reasonable to presume that the local distance metrics would remain approximately constant in small regions.
@@
Remark on Algorithm Choices. In principle, other node embedding and dimensionality reduction methods can be applied to serve our purpose. Our specific choice on HOPE is mainly because of its scalability to large-scale networks and its ability to preserve both local and global structure of the landscape. As for dimensionality reduction, Draganov et al. (2023) has made detailed theoretical comparisons between UMAP and $t$-SNE, and shows that only the normalization significantly impacts their outputs. This thens implies that a majority of the algorithmic differences can be toggled without affecting the embeddings. We choose UMAP here for its better scalability. At the end, the quality of the visualization using our method will continuously grow with the new state-of-the-art in graph represention learning and dimensionality reduction.
Remark on UMAP Hyperparameters. Since HP loss landscapes can vary a lot in terms of dimensionality and total number of configurations, there is no ‘one-size-fits-all’ setup for HPs of UMAP. However, we still provide some general guidelines for HP tuning, which mainly focus on two most important HPs, namely n_neighbors and min_dist. Specifically, n_neighbors controls the balance between local and global structure in the embedding. In general, we found that it is better to set n_neighbors to a value that is larger than the average number of neighbors for each node. For min_dist, which specifies the minimum distance between points in the low-dimensional space, we generally recommend values larger than $0.5$. The reasoning here is that we want the points to be more spread out in the low-dimensional space, and thus prevent the points to be densely distributed in local regions. However, a min_dist that is too large can also cause problems, as different parts of the landscape can get intertwined with each other.
@@ B.2 Landscape Analysis Metrics
0:  A starting configuration $\mathbf{c}$; A neighborhood function $\mathcal{N}_{d}$; A fitness function $f$
1:  while $\mathbf{c}$ is not a local optimum do
2:     $\mathbf{c}^{\prime}_{\text{best}}=\arg\min_{\mathbf{c}^{\prime}\in\mathcal{N}_{1}(\mathbf{c})}f(\mathbf{c}^{\prime})$
3:     if $f(\mathbf{c}^{\prime}_{\text{best}})<f(\mathbf{c})$ then
4:        $\mathbf{c}\leftarrow\mathbf{c}^{\prime}_{\text{best}}$
5:     end if
6:  end while
7:  return $\mathbf{c}$
@@
0:  The set of all configurations $\mathcal{C}$
1:  $\mathcal{V}\leftarrow\emptyset$
2:  $\mathcal{B}\leftarrow\emptyset$
3:  for all $\mathbf{c}\in\mathcal{C}$ do
4:     $\mathbf{c}^{\ell}\leftarrow\textsc{LocalSearch}(\mathbf{c})$
5:     $\mathcal{B}[\mathbf{c}^{\ell}]\leftarrow\mathcal{B}[\mathbf{c}^{\ell}]\cup\{\mathbf{c}\}$
6:  end for
7:  return $\mathcal{V},\mathcal{B}$
@@
Assortativity coefficient. The assortativity coefficient of a network assesses the degree to which nodes tend to be connected to other nodes that are similar w.r.t. some attributes. For example, in a social network, this would mean that people tend to be friends with other people who are similar to themselves in terms of education level, income, race [^4] As one can image, the exact definition of assortativity can depend on whether the target attribute is categorical (i.e., unordered) or numerical (i.e., ordered). Here we focus on the latter as model loss is real-valued., etc. For HP loss landscape, the performance assortativity measures the extent to which HP configurations with similar performance are more likely to be neighbors to each other. Formallly, given a HP loss landscape as a directed graph, and the model loss $\mathcal{L}$ takes values $[L_{1},L_{2},\dots]$, the $\mathcal{L}$-assortativity evaluates the Pearson correlation of the measured loss between pairs of linked configurations and is measured as (Newman, 2003):
-$\mathcal{L}\text{-ast}=\frac{\sum_{i}e_{ii}-\sum_{i}a_{i}b_{i}}{1-\sum_{i}a_{i}b_{i}}$ (2)
@@
where $e_{ij}$ is called mixing matrix entry, which represents the fraction of total edges in the network (i.e., landscape) which connects configurations having performance $\mathcal{L}(\bm{\lambda})=L_{i}$ to configurations having attribute $\mathcal{L}(\bm{\lambda})=L_{j}$. In directed networks like our case, this can be asymmetric, i.e., $e_{ij}\neq e_{ji}$. In addition, $a_{i}=\sum_{j}e_{ij}$ is the portion of edges $(\bm{\lambda}_{u},\bm{\lambda}_{v})$ such that $\mathcal{L}(\bm{\lambda}_{u})=L_{i}$ and $b_{i}=\sum_{j}e_{ji}$ is the portion of edges $(\bm{\lambda}_{v},\bm{\lambda}_{u})$ such that $\mathcal{L}(\bm{\lambda}_{v})=L_{i}$. A high $\mathcal{L}$-ast would imply that configurations with similar performance have strong tendancy to be connected and form local clusters.
@@
Landscape Neutrality. It is often in genetics that a mutation in a single position of a DNA sequence will only lead to negligible change in the expression. Such phenomenon is known as landscape neutrality at a macro level, and for each sequence, this can be quantitatively measured by the mutational robustness (Payne & Wagner, 2019), which is the probability for such non-effective mutation to happen among all its possible mutants. Similar ideas are also applicable to HPO, where altering certain HPs may only result in subtle performance shifts. In particular, we define two neighboring configurations to be neutral if their respective performances differ less than a small fraction $\epsilon$. We choose $\epsilon=0.1\%$ in this paper, since changes below this threshold would make almost no practical meaning. We then define the neutral ratio, denoted as $\nu(\bm{\lambda})$, of a configuration $\bm{\lambda}$ as the portion of its neutral neighbors in its neighborhood. The avarage neutrality of the whole landscape is then defined as:
@@
Neutrality Distance Correlation. While neutrality generally characterizes the expected probability for neutral moves to occur in the whole landscape, it can actually vary across regions. In particular, it is important to investigate whether it is more likely to encounter neutral moves when approaching the global optimum, as in practice, we often find a diminishing gain when tuning towards the best-possible configuration. We quantitatively assess this using the neutrality distance correlation (NDC), which measures the Pearson correlation coefficient between the neutrality of a configuration $\bm{\lambda}$ and its distance to the global optimum, $\mathrm{d}(\bm{\lambda},\bm{\lambda}^{*})$ (equation (4)). We need to note that result derived from this can be potentially affected by the choice of $\epsilon$ for neutrality. To cope with this, we also develop an alternative method for assessing NDC, which is directly based on raw loss differences between configurations. Specifically, for each adaptive walk in the landscape using best-improvement local search (Algorithm 1) that can eventually approach $\bm{\lambda}^{*}$, we measure the Pearson correlation coefficient between $\Delta\mathcal{L}$ for each pair of consecutive configurations $(\bm{\lambda}_{i},\bm{\lambda}_{i-1})$ ($i\geq 2$), and $\mathrm{d}(\bm{\lambda}_{i},\bm{\lambda}^{*})$. We calculate NDC as the average across all such walks. However, in this paper, to keep the consistency with the neutrality metric, we report the results based on the first method.

### v2

@@ 4.3 HP Loss Landscapes with Different Fidelities
Fig. 2 (A-3) and (B-3) show the low-fidelity loss landscapes (denoted as $\mathcal{L}_{\mathrm{test}}^{\mathrm{LF}}$) for CNN and XGBoost with $\alpha=0.1$ training data. From the plots, we observed that $\mathcal{L}_{\mathrm{test}}^{\mathrm{LF}}$ landscapes are highly consistent with $\mathcal{L}_{\mathrm{test}}$ in terms of both structural characteristics and performance rankings. More specifically, as reflected in Fig. 3, all measured FLA metrics of $\mathcal{L}_{\mathrm{test}}^{\mathrm{LF}}$ landscapes exhibited little difference compared to $\mathcal{L}_{\mathrm{test}}$ landscapes across most studied scenarios. For performance rankings, Fig. 4 (B) depicts the distribution of the $3$ similarity indicators between $\mathcal{L}_{\mathrm{test}}^{\mathrm{LF}}$ and $\mathcal{L}_{\mathrm{test}}$ across all datasets for each model. We observed a high Spearman correlation (median $>0.85$) between $\mathcal{L}_{\mathrm{test}}$ and $\mathcal{L}_{\mathrm{test}}^{\mathrm{LF}}$ for most cases, and the $\gamma$-set similarities between the top-$10\%$ configurations are also prominent, with medians larger than $60\%$. These imply that $R(\mathcal{L}_{\mathrm{test}})$ and $R(\mathcal{L}_{\mathrm{test}}^{\mathrm{LF}})$ are highly consistent for the majority of our studied scenarios and there is a large overlap between the promising regions of the two landscapes. In addition, the Shake-up scores yield low values ($\text{median}<0.1$), suggesting that on average the difference between $R(\mathcal{L}_{\mathrm{test}})$ and $R(\mathcal{L}_{\mathrm{test}}^{\mathrm{LF}})$ is less than $10\%$. Note that albeit our general observations, the similarity distributions in Fig. 4 (B) typically have long tails, and there are cases for which the $\mathcal{L}_{\mathrm{test}}^{\mathrm{LF}}$ landscapes largely deviate from the ground truth.
Figure 6: $\rho_{s}$ between $\mathcal{L}_{\mathrm{test}}^{\textrm{LF}}$ and $\mathcal{L}_{\mathrm{test}}$ in NASBench-201 as a function of training epoch.
The NAS landscapes also suggest similar findings. In Fig. 2 (C), we plot the NASBench-101 $\mathcal{L}_{\mathrm{test}}^{\mathrm{LF}}$ landscapes under four different training epochs on CIFAR-10. While it is not surprising that with higher fidelities (i.e., $E$), the $\mathcal{L}_{\mathrm{test}}^{\mathrm{LF}}$ landscapes can better approximate the ground truth $\mathcal{L}_{\mathrm{test}}$ landscape, we noticed that with only $4$ training epochs, the $\mathcal{L}_{\mathrm{test}}^{\mathrm{LF}}$ landscapes can already capture certain prominent regions. We further explored this with NASBench-201, which provides evaluation information after every training epoch. Fig. 6 plots the Spearman correlation between $\mathcal{L}_{\mathrm{test}}^{\mathrm{LF}}$ and $\mathcal{L}_{\mathrm{test}}$ as a function of training epochs $E$. From this we found that with $10$ out of $200$ training epochs, we can already obtain a reasonable ranking.
4.4 HP Loss Landscapes Across Datasets
Figure 7: Importance of top-$5$ HPs (or interactions) as indicated by functional ANOVA method of each model.
Performance rankings. Fig. 2 (A-4) and (B-4) show the $\mathcal{L}_{\mathrm{test}}$ landscapes for CNN and XGBoost on different datasets. We can see that the high-level topography of the $\mathcal{L}_{\mathrm{test}}$ landscape are preserved when transferring to a new task. It is interesting to note that the top regions in the original landscape generally retain their positions, despite the changes in their exact contours. The FLA metrics reported in Fig. 3 also support such observation. In addition, from the similarity metrics reported in Fig. 4, we observed that the measured performance reveal clear Spearman correlations (median $>0.65$) across datasets. In practice, stakeholders are more concerned about the top performing regions in the landscape. For this, the overlap between the top-$10\%$ regions, as indicated by the $\gamma$-set similarity, also achieves medians around $40\%$. We also found that the landscape similarities for tabular tasks are correlated with simple dataset meta-features (see Fig. 15 in Appendix F), and there have been works that study such features in more depth (e.g., [79]).
HP importance and interaction. In addition to performance rankings, Fig. 7 illustrates the contribution of each HP and their interactions to model performance assessed by the functional ANOVA method. The results indicate that some (or combination terms of) HPs are typically important for many datasets for a given model. For example, the depth-conv contributes a large portion of variance to model performance for CNN, and so does its interactions with width-conv. Such knowledge can also be utilized to expedite the HPO process.
Our additional results for NAS-Bench-201 across CIFAR-10/100 as well as ImageNet-16-120 datasets also revealed similar highly-transferable conclusions (see Fig. 2 D1-D3). However, again, we note that our findings are based on general trends, and there are outliers in Fig. 4 that reveal little landscape similarity between tasks. Also, the results on NASBench-360 (see Fig. 2 D4-D5), which evaluates the search space of NASBench-201 on two tasks very different from classic image classification datasets, exhibited considerable deviations from the landscapes in Fig. 2 D1-D3. Considering that these two tasks are rooted in biomechanics and computational physics respectively, the resulting difference might be reasonable. However, this evokes the need for including a more diverse set of tasks to further validate our findings, and we leave this for future works.
5 Discussions and Conclusions
Many studies have significantly advanced the methodological developments of HPO. Yet, the understanding of the underlying HP loss landscapes of ML models as well as their correlations, which are the foundations of HPO, has been largely overlooked. By analyzing $1,500+$ HP loss landscapes of $6$ representative ML models induced on different scenarios with fine-grained visualizations and dedicated FLA metrics, we found a considerable resemblance between test loss landscapes and their training counterparts, albeit the choice of certain HP combinations and datasets can lead to significant discrepancies. We also showed that the topographies of HP loss landscapes with lower fidelities, in general, are highly consistent with the full-fidelity landscapes, and the same applies to landscapes of the same model induced on different datasets. We further revealed a unified portrait of the HP loss landscapes across our studied models in terms of their structural characteristics. For example, they tend to be fairly smooth, nearly unimodal, and featured with a high degree of neutrality.
Our results in part reaffirm isolated patterns observed in prior works [31, 32, 33, 34, 35] regarding landscape modality and neutrality, and extend them to broader scenarios. At the same time, we provide the first landscape-level evidence to support the effectiveness of transfer learning and multi-fidelity methods, which previously have been mainly based on rough intuition and empirical performance. Moreover, the high-level structural similarities shared by our studied HP loss landscapes induced on different models assist in explaining why many HPO solvers often work robustly across different systems. Also, given the smooth topography of these landscapes, we also highlight the importance of elaborated search space design, including both the selection of HPs and the choice of value domains.
@@ 5 Discussions and Conclusions
Despite these caveats, our exploratory study may provide a useful perspective to rethink HPO algorithms by considering the topography of the underlying HP loss landscapes upon which they traverse. There could also be nuanced scrutiny of the scenarios discussed in this study, e.g., the relationship between dataset meta-features and the landscape characteristics. Also, our developed FLA framework could serve as a useful lens for future studies on a broader range of search landscapes in the ML community (e.g., NAS, AutoRL, and AutoML), and beyond (e.g., protein fitness landscapes [?] toward AI for science). We hope our analysis opens a new avenue towards understanding the fitness landscapes of black-box systems and explainable AI in optimization.
Appendix A Summary of Computational Resrouces Used
[…]
Appendix B Details of the FLA Framework
@@ B.1 Landscape Visualization Method
HOPE node embedding. To preserve the intrinsic neighborhood relationship of HP configurations, our proposed landscape visualization method first needs to learn a low-dimensional embedding for each configuration in the landscape (i.e., each node in the graph). In this paper, we choose HOPE [^1] We use the implementation in Karateclub package. node embedding method to serve this purpose, as it is able to capture asymmetric high-order proximity in directed networks. Specifically, for a directed network [^2] Most manipulations and analyses of HP loss landscapes as complex networks are grounded on NetworkX and Pandas package., if there is a directed edge from vertex $v_{i}$ to vertex $v_{j}$ and from vertex $v_{j}$ to vertex $v_{k}$, it is more likely to have a edge from $v_{i}$ to $v_{k}$, but not from $v_{k}$ to $v_{i}$. In order to preserve such asymmetric transitivity, HOPE learns two vertex embedding vectors $U^{s},U^{t}\in\mathbf{R}^{|V|\times d}$, which is called source and target embedding vectors, respectively. After constructing the high-order proximity matrix $S$ from $4$ proximity measures, i.e., Katz Index, Rooted PageRank, Common Neighbors and AdamicAdar. HOPE finally learns vertex embeddings by solving the a matrix factorization problem:
@@
UMAP dimensionality reduction. While HOPE (as well as any other node embedding methods) could generate vectorized embeddings for configurations, these embeddings typically are still high-dimensional and not compatible for $2$D visualization. In light of this, we further apply UMAP [^3] We use the implementation in UMAP package. to project the HOPE embeddings into a $2$D space, which is based on three assumptions, and here we provide brief discussion on whether they hold for our case:
[…]
    Data is uniformly distributed on Riemannian manifold. As the distribution of HOPE embeddings essentially depends on the connectivity pattern of the graph, for HP loss landscapes, we can expect this assumption to hold. This is because the number of neighbors for each node in the graph would largely stay the same based on our neighborhood definition. Therefore, there will be no significance variation in density across the graph, and the HOPE embeddings in turn should approximately have a uniform distribution.
[…]
    The Riemannian metric is locally constant. The distances between HOPE embeddings correlate directly with the local connectivity patterns within the network. Given the stability in neighborhood structures as we discussed above, it is reasonable to presume that local distance metrics would remain approximately constant in small regions.
[…]
@@
Remark on algorithmic choices. In principle, other node embedding and dimensionality reduction methods can be applied to serve our purpose. Our specific choice on HOPE is mainly because of its scalability to large-scale networks and its ability to preserve both local and global structure of the landscape. As for dimensionality reduction, [83] has made detailed theoretical comparisons between UMAP and $t$-SNE, and shows that only the normalization significantly impacts their outputs. This then implies that a majority of the algorithmic differences can be toggled without affecting the embeddings. We choose UMAP here for its better scalability. At the end, the quality of the visualization using our method will continuously grow with the new state-of-the-art in graph represention learning and dimensionality reduction.
Remark on UMAP hyperparameters. Since HP loss landscapes can vary a lot in terms of dimensionality and total number of configurations, there is no ‘one-size-fits-all’ setup for HPs of UMAP. Yet, here we provide some empirical guidelines for HP tuning, which mainly focus on two most important HPs, namely n_neighbors and min_dist. Specifically, n_neighbors controls the balance between local and global structure in the embedding. In general, we found that it is better to set n_neighbors to a value that is larger than the average number of neighbors for each node. For min_dist, which specifies the minimum distance between points in the low-dimensional space, we recommend values larger than $0.5$. The reasoning here is that we want the points to be more spread out in the low-dimensional space, and thus prevent the points to be densely distributed in local regions. However, a min_dist that is too large can also cause problems, as different parts of the landscape can get intertwined with each other.
@@ B.2 Landscape Analysis Metrics
0:  A starting configuration $\mathbf{c}$; A neighborhood function $\mathcal{N}_{d}$; A fitness function $f$ 1:  while $\mathbf{c}$ is not a local optimum do 2:    $\mathbf{c}^{\prime}_{\text{best}}=\arg\min_{\mathbf{c}^{\prime}\in\mathcal{N}_{1}(\mathbf{c})}f(\mathbf{c}^{\prime})$ 3:    if $f(\mathbf{c}^{\prime}_{\text{best}})<f(\mathbf{c})$ then 4:       $\mathbf{c}\leftarrow\mathbf{c}^{\prime}_{\text{best}}$ 5:    end if 6:  end while 7:  return $\mathbf{c}$
@@
0:  The set of all configurations $\mathcal{C}$ 1:  $\mathcal{V}\leftarrow\emptyset$ 2:  $\mathcal{B}\leftarrow\emptyset$ 3:  for all $\mathbf{c}\in\mathcal{C}$ do 4:    $\mathbf{c}^{\ell}\leftarrow\textsc{LocalSearch}(\mathbf{c})$ 5:    $\mathcal{B}[\mathbf{c}^{\ell}]\leftarrow\mathcal{B}[\mathbf{c}^{\ell}]\cup\{\mathbf{c}\}$ 6:  end for 7:  return $\mathcal{V},\mathcal{B}$
@@
Assortativity coefficient. The assortativity coefficient of a network assesses the degree to which nodes tend to be connected to other nodes that are similar w.r.t. some attributes. For example, in a social network, this would mean that people tend to be friends with other people who are similar to themselves in terms of education level, income, race [^4] As one can image, the exact definition of assortativity can depend on whether the target attribute is categorical (i.e., unordered) or numerical (i.e., ordered). Here we focus on the latter as model loss is real-valued., etc. For HP loss landscape, the performance assortativity measures the extent to which HP configurations with similar performance are more likely to be neighbors to each other. Formallly, given a HP loss landscape as a directed graph, and the model loss $\mathcal{L}$ takes values $[L_{1},L_{2},\dots]$, the $\mathcal{L}$-assortativity evaluates the Pearson correlation of the measured loss between pairs of linked configurations and is measured as [58]:
$\mathcal{L}_{\text{ast}}=\frac{\sum_{i}e_{ii}-\sum_{i}a_{i}b_{i}}{1-\sum_{i}a_{i}b_{i}}$ (2)
@@
where $e_{ij}$ is called mixing matrix entry, which represents the fraction of total edges in the network (i.e., landscape) that connects configurations having performance $\mathcal{L}(\bm{\lambda})=L_{i}$ to configurations having attribute $\mathcal{L}(\bm{\lambda})=L_{j}$. In directed networks like our case, this can be asymmetric, i.e., $e_{ij}\neq e_{ji}$. In addition, $a_{i}=\sum_{j}e_{ij}$ is the portion of edges $(\bm{\lambda}_{u},\bm{\lambda}_{v})$ such that $\mathcal{L}(\bm{\lambda}_{u})=L_{i}$ and $b_{i}=\sum_{j}e_{ji}$ is the portion of edges $(\bm{\lambda}_{v},\bm{\lambda}_{u})$ such that $\mathcal{L}(\bm{\lambda}_{v})=L_{i}$. A high $\mathcal{L}_{\text{ast}}$ would imply that configurations with similar performance have strong tendancy to be connected and form local clusters.
@@
Landscape neutrality. It is often in genetics that a mutation in a single position of a DNA sequence will only lead to negligible change in the expression. Such phenomenon is known as landscape neutrality at a macro level, and for each sequence, this can be quantitatively measured by the mutational robustness [84], which is the probability for such non-effective mutation to happen among all its possible mutants. Similar ideas are also applicable to HPO, where altering certain HPs may only result in subtle performance shifts. In particular, we define two neighboring configurations to be neutral if their respective performances differ less than a small fraction $\epsilon$. We choose $\epsilon=0.1\%$ in this paper, since changes below this threshold would make almost no practical meaning. We then define the neutral ratio, denoted as $\nu(\bm{\lambda})$, of a configuration $\bm{\lambda}$ as the portion of its neutral neighbors in its neighborhood. The avarage neutrality of the whole landscape is then defined as:
@@
Neutrality distance correlation. While neutrality generally characterizes the expected probability for neutral moves to occur in the whole landscape, it can actually vary across regions. In particular, it is important to investigate whether it is more likely to encounter neutral moves when approaching the global optimum, as in practice, we often find a diminishing gain when tuning towards the best-possible configuration. We quantitatively assess this using the neutrality distance correlation (NDC), which measures the Pearson correlation coefficient between the neutrality of a configuration $\bm{\lambda}$ and its distance to the global optimum, $\mathrm{d}(\bm{\lambda},\bm{\lambda}^{*})$ (equation (4)). Specifically, for each adaptive walk in the landscape using best-improvement local search (Algorithm 1) that can eventually approach $\bm{\lambda}^{*}$, we measure the Pearson correlation coefficient between $\Delta\mathcal{L}$ for each pair of consecutive configurations $(\bm{\lambda}_{i},\bm{\lambda}_{i-1})$ ($i\geq 2$), and $\mathrm{d}(\bm{\lambda}_{i},\bm{\lambda}^{*})$. We calculate NDC as the average across all such walks.